The diameter of a metallic ball is 4.2 cm. What is the mass of the ball, if the density of the metal is 8.9 g per cm3? •

Grade:12th pass

2 Answers

Pawan Prajapati
3 years ago
Diameter of the ball = 4.2 cm. Radius of the ball = 2.1 cm. 4 2'Z Volume of the ball = x 7x (2.1>3 cm = 38.808 cm. Mass of the ball = density x volume ' = 8.9 x 38.808 g = 345.39 g (approx.)
SUVAN SETHY
3 years ago
Diameter = 4.2 cm  , Density = 8.9g/ccRadius = 2.1cm  , volume =  4/3 *22/7*2.1
Mass =  Density* Volume =>  Mass =  8.9 * 4/3 * 22/7 * 2.13 =  345.3912 g = 0.345 kg